You are creating a portfolio of Stock D and Stock BW (from earlier).

You are

investing $2,000 in Stock BW and $3,000 in Stock D.

The expected return and

standard deviation of Stock D is 8% and 10.65% respectively.

The correlation

coefficient between BW and D is 0.75.What is the expected return and standard

deviation of the portfolio?



1
Expert's answer
2021-05-07T09:54:45-0400

Assuming the ex pected return and standard deviation of stock BW is 9% and 13.15% respectively;


(i) Determining the portfolio expected return


"W_{BW}=\frac{\$2,000}{\$5,000}=0.4"


"W_D=\frac{\$3,000}{\$5,000}=0.6"


"R_P=(W_{BW})(R_{BW})+(W_D)(R_D)"


"=(0.4)(9\%)+(0.6)(8\%)"


"=8.4\%"


(ii)

Determining portfolio standard deviation

Two-asset portfolio:




determining the portfolio standard deviation

"S_p=\sqrt{0.0028+0.0025+0.0025+0.0041}"


"=\sqrt{0.0119}"


"=0.1091"